# An In-depth Evaluation of GPT-4 in Sentence Simplification with Error-based Human Assessment

### v1

@@ 4.2. Replicated Control-T5
We replicated the Control-T5 model (Sheang and Saggion, 2021). We started by fine-tuning the T5-base model (Raffel et al., 2020) with the WikiLarge dataset and then evaluated it on the ASSET and Turk’s test sets. Unlike the original study, which did not train on or evaluate on Newsela, we incorporated this dataset. We employed Optuna (Akiba et al., 2019) for hyperparameter optimization, a method consistent with the approach used in the original study with the WikiLarge dataset. This optimization process focused on adjusting the batch size, the number of epochs, the learning rate, and the control token ratios. We refer the reader to Appendix A.2 for the optimal model configuration we achieved.
@@ 5. Human Evaluation
Automatic metrics provide a fast and cost-effective way for evaluating simplification but struggle to cover all the aspects; they are designed to capture only specific aspects such as the similarity between the output and a reference. Furthermore, the effectiveness of some automatic metrics has been challenged in previous studies (Sulem et al., 2018a; Alva-Manchego et al., 2021). Human evaluation, which is often viewed as the gold standard evaluation, may be a more reliable method to determine the quality of simplification. As we discuss in detail in the following section, achieving a balance between interpretability and consistency among annotators is a challenge in sentence simplification. To address this challenge, we have crafted an error-based approach and made efforts in the annotation process such as mandating discussions among annotators to achieve consensus and implementing strict checks to ensure the quality of assessments.
@@ 5.1.1. Challenge in Current Human Evaluation
Sentence simplification is expected to make the original sentence simpler while maintaining grammatical integrity and not losing important information. A common human assessment approach involves rating sentence-level automatic simplification outputs by comparing them to source sentences in three aspects: fluency, meaning preservation, and simplicity (Kriz et al., 2019; Jiang et al., 2020; Alva-Manchego et al., 2021; Maddela et al., 2021). However, sentence simplification involves various transformations, such as paraphrasing, deletion, and splitting, which affect both the lexical and structural aspects of a sentence. Sentence-level scores are difficult to interpret; they do not clearly indicate whether the transformations simplify or complicate the original sentence, maintain or alter the original meaning, or are necessary or unnecessary. Therefore, such evaluation approach falls short in comprehensively assessing the models’ capabilities.
This inadequacy has led to a demand for more detailed and nuanced human assessment methods. Recently, the SALSA framework, introduced by Heineman et al. (Heineman et al., 2023) aimed to provide clearer insights through comprehensive human evaluation, consider both the successes and failures of a simplification system. This framework categorizes transformations into $21$ linguistically-grounded edit types across conceptual, syntactic, and lexical dimensions to facilitate detailed evaluation. However, due to the detailed categorization, it faces challenges in ensuring consistent interpretations across annotators. This inconsistency frequently leads to low inter-annotator agreement, thereby undermining the reliability of the evaluation. We argue that such extensive and fine-grained classifications are difficult for annotators to understand, particularly those without a linguistic background, making it challenging for them to keep consistency.
@@ 5.1.2. Error-based Human Evaluation
To overcome the trade-off between interpretability and consistency in evaluations, we design our error-based human evaluation framework. Our approach focuses on identifying and evaluating key failures generated by advanced LLMs in important aspects of sentence simplification. We aim to cover a broad range of potential failures while making the classification easy for annotators. Our approach reduces the categories to seven types while ensuring comprehensive coverage of common failures. In the study on sentence simplification evaluation of LLMs conducted by Kew et al. (Kew et al., 2023), while the annotation of common failures is also incorporated, it is noteworthy that the types of failures addressed are very limited and they selected only a handful of output samples for annotation.
While not intended for LLM-based simplification, a few previous studies have incorporated error analysis to assess their sequence-to-sequence simplification models (Kriz et al., 2019; Cooper and Shardlow, 2020; Maddela et al., 2021). Starting from the error types established in these studies, we included ones that might also applicable in the outputs of advanced LLMs. Specifically, we conducted a preliminary investigation on ChatGPT-3.5 simplification outputs on the ASSET dataset. [^4] At the time of this investigation, GPT-4 was not publicly available. As a result, we adopted errors of Altered Meaning, issues with Coreference, Repetition, and Hallucination, while omitted errors deemed unlikely, such as the ungrammatical error. Additionally, we identified a new category of error based on our investigation: Lack of Simplicity. We observed that ChatGPT-3.5 often opted for more complex expressions rather than simpler ones, which is counterproductive for sentence simplification. Recognizing this as a significant issue, we included it in our error types. We also refined the categories for altered meaning and lack of simplicity by looking into the specific types of changes they involve. Instead of listing numerous transformations like the SALSA framework (Heineman et al., 2023), we classified these transformations into two simple categories based on their effects on the source sentence: lexical and structural changes. This categorization leads to four error types: Lack of Simplicity-Lexical, Lack of Simplicity-Structural, Altered Meaning-Lexical, and Altered Meaning-Structural.
@@
Lack of Simplicity Lexical
The simplified sentence uses more intricate lexical expression(s) to replace part(s) of the original sentence.
For Rowling, this scene is important because it shows Harry’s bravery…
Rowling considers the scene significant because it portrays Harry’s courage…
Structural
The simplified sentence modifies the grammatical structure, and it increases the difficulty of reading.
The other incorporated cities on the Palos Verdes Peninsula include…
Other cities on the Palos Verdes Peninsula include…, which are also incorporated.
Altered Meaning Lexical
Significant deviation in the meaning of the original sentence due to lexical substitution(s).
The Britannica was primarily a Scottish enterprise.
The Britannica was mainly a Scottish endeavor.
Structural
Significant deviation in the meaning of the original sentence due to structural changes.
Gimnasia hired first famed Colombian trainer Francisco Maturana, and then Julio César Falcioni.
Gimnasia hired two famous Colombian trainers, Francisco Maturana and Julio César Falcioni.
Coreference
A named entity critical to understanding the main idea is replaced with a pronoun or a vague description.
Sea slugs dubbed sacoglossans are some of the most…
These are some of the most…
Repetition
Unnecessary duplication of sentence fragments
The report emphasizes the importance of sustainable practices.
The report emphasizes the importance, the significance, and the necessity of sustainable practices.
Hallucination
Inclusion of incorrect or unrelated information not present in the original sentence.
In a short video promoting the charity Equality Now, Joss Whedon confirmed that ”Fray is not done, Fray is coming back.
Joss Whedon confirmed in a short promotional video for the charity Equality Now that Fray will return, although the story is not yet finished.
@@ 5.2.1. Task $1$: Error Identification
Task $1$ follows our error-based human evaluation detailed in Section 5.1. We sampled $300$ source sentences from Turk, ASSET, and Newsela test sets, along with simplification outputs generated by GPT-4 and Control-T5, resulting in a total of $1800$ complex-simple sentence pairs. The annotation period spanned from October $2023$ to February $2024$.
Annotators were instructed to identify and label errors within each sentence pair according to predefined guidelines. To overcome the trade-off between detailed granularity and annotator agreement, all annotators involved in this task participated in discussion sessions led by one of the authors. These sessions required annotators to share their individual labelings, which were then collectively reviewed during discussions until reaching the consensus. There were eight discussion sessions, each lasting approximately three hours, for a total of $24$ hours.
@@
As annotators, we used second-language learners with advanced English proficiency expecting that they are more sensitive to variations in textual difficulty based on their language-learning experiences. In addition, given that second language learners stand to benefit significantly from sentence simplification applications, involving them as evaluators seems most appropriate. All of our annotators were graduate students associated with our organization. The compensation rate for this task was set at ¥$100$ JPY (approximately \$$0.67$ USD) per sentence pair. For quality control, annotators had to pass a qualification test before participating in the task. This qualification test comprises annotation guidelines and four complex-simple sentence pairs. Each pair contains various errors predefined by the author. All submissions to this test were manually reviewed. Three annotators were selected for this task based on their high accuracy in identifying errors, including specifying the error type, location, and rationale.
@@ 5.2.2. Task $2$: Likert Scale Rating
Following the convention of previous studies, we also include the rating approach on fluency, meaning preservation, and simplicity using a 1 to 3 Likert scale as Task $2$. In this task, annotators evaluate all simplification outputs generated by GPT-4 and Control-T5 across three test sets, by comparing them with their corresponding source sentences. In particular, for the Newsela dataset, reference simplifications from the test set were also included. We assume that models trained or tuned on this dataset which is characterized by deletion, may produce shorter outputs, potentially impacting meaning preservation scores. To ensure fairness, we compare human evaluation of model-generated simplifications against that of Newsela reference simplifications for a more objective evaluation. The evaluation in Task $2$ covered a total of $4667$ complex-simple sentence pairs. The annotation period spanned from October $2023$ to February $2024$.
To address the challenge of annotator consistency, we implemented specific guidelines during the annotation phase. Annotators were advised to avoid neutral positions (‘2’ on our scale) unless faced with genuinely challenging decisions. This approach encouraged a tendency towards binary choices, i.e., ‘1’ for simplification outputs that are disfluent, lose a lot of original meaning, or are not simpler, and ‘3’ for simplification outputs that are fluent, preserve meaning, and are much simpler. To ensure quality, one of the authors reviewed $200$ pairs of sampled submissions from each annotator. If any issues were identified, such as an annotator rate inconsistently for sentence pairs with similar problems, they were required to revise and resubmit their annotations.
@@
Same with Task $1$, we employed second-language learners with advanced English proficiency, who were graduate students associated with our organization. Annotator candidates had to pass a qualification test before participating in the task. This qualification test comprises annotation guidelines and five complex-simple sentence pairs. Candidates were instructed to rate fluency, meaning preservation, and simplicity on each simplification output. Three annotators were selected based on their high inter-annotator agreement, demonstrated by the Intraclass Correlation Coefficient (ICC) (Shrout and Fleiss, 1979) score of $0.72$, indicating a substantial agreement. The compensation rate for this task was set at ¥$40$ JPY (approximately \$$0.27$ USD) per sentence pair.
@@
We assessed inner-annotator agreement through the overlapping rate of ratings across three annotators, as detailed in Table 4. The overlapping rate was calculated by the proportion of identical ratings for a given simplification output. [^5] We also tried Fleiss’ kappa, Krippendorff’s alpha, and ICC; however, they resulted in degenerate scores due to too-high agreements on mostly binary judgments. In the fluency dimension, both models demonstrate strong agreement, with overlapping rates between $96.9\%$ and $99.6\%$. In meaning preservation and simplicity, these dimensions exhibit comparably more variability in ratings, with a broader range of agreement. Annotators found it more subjective to assess meaning preservation and simplicity, as these aspects required direct comparison with the source sentences. Nevertheless, mid to high agreement levels were still achieved, showing the consistency of our annotation.
@@
Turk ASSET Newsela
Dimension GPT-4 T5 GPT-4 T5 GPT-4 T5 Reference
Fluency $98.1$ $97.8$ $97.8$ $96.9$ $99.6$ $98.4$ $98.0$
Meaning Preservation $86.4$ $57.9$ $84.7$ $58.2$ $66.6$ $87.3$ $62.6$
Simplicity $77.7$ $85.5$ $68.2$ $90.3$ $81.3$ $95.5$ $76.5$
@@ 6. Annotation Result Analysis
Our comprehensive analysis of annotation data revealed that, overall, GPT-4 generates fewer erroneous simplification outputs compared to Control-T5, demonstrating higher ability in simplification. GPT-4 excels in maintaining the original meaning, whereas Control-T5 often falls short in this dimension. Nevertheless, GPT-4 is not without flaws; its most frequent mistake involves substituting lexical expressions with more complex ones.
@@ 6.1. Analysis of Task 1: Error Identification
This section presents a comparative analysis of erroneous simplification outputs generated by GPT-4 and Control-T5, focusing on error quantification and type analysis. We assessed erroneous simplification outputs across three datasets: Turk, ASSET, and Newsela, defining an erroneous output as one containing at least one error. Overall, Table 5 shows that GPT-4 with fewer errors than Control-T5 across all datasets. This performance difference underscores GPT-4’s superior performance in simplification tasks. In the following, we reveal characteristics of errors in GPT-4 and Control-T5 simplification outputs.
Table 5. Comparison of the number of erroneous simplification outputs generated by GPT-4 and Control-T5
Test set GPT-4 Control-T5
Turk ($300$ samples) $45$ $114$
ASSET ($300$ samples) $64$ $100$
Newsela ($300$ samples) $73$ $115$
Total ($900$ samples) $182$ $329$

### v2

@@ 4.2. Replicated Control-T5
We replicated the Control-T5 model (Sheang and Saggion, 2021). We started by fine-tuning the T5-base model (Raffel et al., 2020) with the WikiLarge dataset and then evaluated it on the ASSET and Turk’s test sets. Unlike the original study, which did not train on or evaluate on Newsela, we incorporated this dataset. We employed Optuna (Akiba et al., 2019) for hyperparameter optimization, a method consistent with the approach used in the original study with the WikiLarge dataset. This optimization process focused on adjusting the batch size, the number of epochs, the learning rate, and the control token ratios. Note that we did not evaluate Control-T5’s performance on SimPA since the training dataset is not available. We refer the reader to Appendix A.2 for the optimal model configuration we achieved.
@@ 5. Human Evaluation
Automatic metrics provide a fast and cost-effective way for evaluating simplification but struggle to cover all the aspects; they are designed to capture only specific aspects such as the similarity between the output and a reference. Furthermore, the effectiveness of some automatic metrics has been challenged in previous studies (Sulem et al., 2018a; Alva-Manchego et al., 2021). Human evaluation, which is often viewed as the gold standard evaluation, may be a more reliable method to determine the quality of simplification. As we discuss in detail in the following section, achieving a balance between interpretability and consistency among annotators is a challenge in sentence simplification. To address this challenge, we have crafted an error-based approach and made efforts in the annotation process, such as mandating discussions among annotators to achieve consensus and implementing strict checks to ensure the quality of assessments.
@@ 5.1.1. Challenge in Current Human Evaluation
Sentence simplification is expected to make the original sentence simpler while maintaining grammatical integrity and not losing important information. A common human assessment approach involves rating sentence-level automatic simplification outputs by comparing them to source sentences in three aspects: fluency, meaning preservation, and simplicity (Kriz et al., 2019; Jiang et al., 2020; Alva-Manchego et al., 2021; Maddela et al., 2021). However, sentence simplification involves various transformations, such as paraphrasing, deletion, and splitting, which affect both the lexical and structural aspects of a sentence. Sentence-level scores are difficult to interpret; they do not indicate whether the transformations simplify or complicate the original sentence, maintain or alter the original meaning, or are necessary or unnecessary. Therefore, such evaluation approach falls short in comprehensively assessing the models’ capabilities.
This inadequacy has led to a demand for more detailed and nuanced human assessment methods. Recently, the SALSA framework, introduced by Heineman et al. (Heineman et al., 2023), aimed to provide clearer insights through comprehensive human evaluation and consider both the successes and failures of a simplification system. This framework categorizes transformations into $21$ linguistically-grounded edit types across conceptual, syntactic, and lexical dimensions to facilitate detailed evaluation. However, due to the detailed categorization, it faces challenges in ensuring consistent interpretations across annotators. This inconsistency frequently leads to low inter-annotator agreement, thereby undermining the reliability of the evaluation. We argue that such extensive and fine-grained classifications are difficult for annotators to understand, particularly those without a linguistic background, making it challenging for them to maintain consistency.
@@ 5.1.2. Error-based Human Evaluation
To overcome the trade-off between interpretability and consistency in evaluations, we design our error-based human evaluation framework. Our approach focuses on identifying and evaluating key failures generated by advanced LLMs in important aspects of sentence simplification. We aim to cover a broad range of potential failures while making the classification easy for annotators. Our approach reduces the categories to seven types while ensuring comprehensive coverage of common failures. In the study on sentence simplification evaluation of LLMs conducted by Kew et al. (Kew et al., 2023), while the annotation of common failures is also incorporated, it is noteworthy that the types of failures addressed were very limited, and they selected only a handful of output samples for annotation.
While not intended for LLM-based simplification, a few previous studies have incorporated error analysis to assess their sequence-to-sequence simplification models (Kriz et al., 2019; Cooper and Shardlow, 2020; Maddela et al., 2021). Starting from the error types established in these studies, we included ones that might also be applicable in the outputs of advanced LLMs. Specifically, we conducted a preliminary investigation of ChatGPT-3.5 simplification outputs on the ASSET dataset. [^5] At the time of this investigation, GPT-4 was not publicly available. As a result, we adopted errors of Altered Meaning, issues with Coreference, Repetition, and Hallucination, while omitted errors deemed unlikely, such as the ungrammatical error. Additionally, we identified a new category of error based on our investigation: Lack of Simplicity. We observed that ChatGPT-3.5 often opted for more complex expressions rather than simpler ones, which is counterproductive for sentence simplification. Recognizing this as a significant issue, we included it in our error types. We also refined the categories for altered meaning and lack of simplicity by looking into the specific types of changes they involve. Instead of listing numerous transformations like the SALSA framework (Heineman et al., 2023), we classified these transformations into two simple categories based on their effects on the source sentence: lexical and structural changes. This categorization leads to four error types: Lack of Simplicity-Lexical, Lack of Simplicity-Structural, Altered Meaning-Lexical, and Altered Meaning-Structural.
@@
Lack of Simplicity Lexical The simplified sentence uses more intricate lexical expression(s) to replace part(s) of the original sentence. For Rowling, this scene is important because it shows Harry’s bravery… Rowling considers the scene significant because it portrays Harry’s courage…
Structural The simplified sentence modifies the grammatical structure, and it increases the difficulty of reading. The other incorporated cities on the Palos Verdes Peninsula include… Other cities on the Palos Verdes Peninsula include…, which are also incorporated.
Altered Meaning Lexical Significant deviation in the meaning of the original sentence due to lexical substitution(s). The Britannica was primarily a Scottish enterprise. The Britannica was mainly a Scottish endeavor.
Structural Significant deviation in the meaning of the original sentence due to structural changes. Gimnasia hired first famed Colombian trainer Francisco Maturana, and then Julio César Falcioni. Gimnasia hired two famous Colombian trainers, Francisco Maturana and Julio César Falcioni.
Coreference A named entity critical to understanding the main idea is replaced with a pronoun or a vague description. Sea slugs dubbed sacoglossans are some of the most… These are some of the most…
Repetition Unnecessary duplication of sentence fragments The report emphasizes the importance of sustainable practices. The report emphasizes the importance, the significance, and the necessity of sustainable practices.
Hallucination Inclusion of incorrect or unrelated information not present in the original sentence. In a short video promoting the charity Equality Now, Joss Whedon confirmed that ”Fray is not done, Fray is coming back. Joss Whedon confirmed in a short promotional video for the charity Equality Now that Fray will return, although the story is not yet finished.
@@ 5.2.1. Task $1$: Error Identification
Task $1$ follows our error-based human evaluation detailed in Section 5.1. We sampled $300$ source sentences from Turk, ASSET, and Newsela test sets, along with simplification outputs generated by GPT-4, Qwen2.5-72B, Llama-3.2-3B, and Control-T5, resulting in a total of $3,600$ complex-simple sentence pairs. Additionally, we sampled $300$ source sentences from SimPA and generated simplification outputs using the three LLMs, contributing an additional $900$ complex-simple sentence pairs. Altogether, this resulted in $4500$ complex-simple sentence pairs. The annotation process was conducted over two periods: from October $2023$ to February $2024$, and from October $2024$ to December $2024$.
Annotators were instructed to identify and label errors within each sentence pair according to predefined guidelines. To overcome the trade-off between detailed granularity and annotator agreement, all annotators involved in this task participated in discussion sessions led by one of the authors. These sessions required annotators to share their individual labelings, which were then collectively reviewed during discussions until a consensus was reached. There were $20$ discussion sessions, each lasting approximately three hours, for a total of $60$ hours.
@@
As annotators, we used second-language learners with advanced English proficiency, expecting that they would be more sensitive to the fine-grained level of variations in textual difficulty based on their language-learning experiences. In addition, given that second language learners stand to benefit significantly from sentence simplification applications, involving them as evaluators seems most appropriate. All of our annotators were graduate students or alumni associated with our organization. The compensation rate for this task was set at ¥$100$ JPY (approximately \$$0.67$ USD) per sentence pair. For quality control, annotators had to pass a qualification test before participating in the task. This qualification test comprises annotation guidelines and four complex-simple sentence pairs. Each pair contains various errors predefined by the author. All submissions to this test were manually reviewed. Seven annotators were selected for this task based on their high accuracy in identifying errors, including specifying the error type, location, and rationale. They come from Brazil, China, Italy, Indonesia, and Israel.
@@ 5.2.2. Task $2$: Likert Scale Rating
Following the convention of previous studies, we also include the rating approach on fluency, meaning preservation, and simplicity using a 1 to 3 Likert scale as Task $2$. In this task, annotators evaluate all simplification outputs generated by LLMs and Control-T5 across four test sets, by comparing them with their corresponding source sentences. In particular, for the Newsela dataset, reference simplifications from the test set were also included. We assume that models trained or tuned on this dataset, which is characterized by deletion, may produce shorter outputs, potentially impacting meaning preservation scores. To ensure fairness, we compare the human evaluation of model-generated simplifications against that of Newsela reference simplifications for a more objective evaluation. The evaluation in Task $2$ covered a total of 11,548 complex-simple sentence pairs. The annotation process was conducted over two periods: from October $2023$ to February $2024$, and from October $2024$ to December $2024$.
To address the challenge of annotator consistency, we implemented specific guidelines during the annotation phase. Annotators were advised to avoid neutral positions (‘2’ on our scale) unless faced with genuinely challenging decisions. This approach encouraged a tendency towards binary choices, i.e., ‘1’ for simplification outputs that are disfluent, lose a lot of original meaning, or are not simpler, and ‘3’ for simplification outputs that are fluent, preserve meaning, and are much simpler. To ensure quality, one of the authors reviewed $200$ pairs of sampled submissions from each annotator. If any issues were identified, such as an annotator rating inconsistently for sentence pairs with similar problems, they were required to revise and resubmit their annotations.
@@
Same with Task $1$, we used second-language learners with advanced English proficiency, who were graduate students or alumni associated with our organization. Additionally, we included a native speaker with English education experience, as this candidate’s evaluations demonstrated comparable reliability in overall assessment quality. Annotator candidates had to pass a qualification test before participating in the task. This qualification test comprises annotation guidelines and five complex-simple sentence pairs. Candidates were instructed to rate fluency, meaning preservation, and simplicity on each simplification output. Seven annotators were selected based on their high inter-annotator agreement, demonstrated by the Intraclass Correlation Coefficient (ICC) (Shrout and Fleiss, 1979) score of $0.62$, indicating a substantial agreement. They come from Brazil, China, Italy, Indonesia, Malaysia, and the United States. The compensation rate for this task was set at ¥$40$ JPY (approximately \$$0.27$ USD) per sentence pair.
@@
We assess inner-annotator agreement through the overlapping rate of ratings across three annotators, as detailed in Table 4. The overlapping rate is calculated by the proportion of identical ratings for a given simplification output. [^6] We also tried Fleiss’ kappa, Krippendorff’s alpha, and ICC; however, they resulted in degenerate scores due to too-high agreements on mostly binary judgments. In the fluency dimension, all models demonstrate strong agreement, with overlapping rates between $96.9\%$ and $99.7\%$. In meaning preservation and simplicity, these dimensions exhibit comparably more variability in ratings, with a broader range of agreement. Annotators found it more subjective to assess meaning preservation and simplicity, as these aspects required direct comparison with the source sentences. Nevertheless, mid to high agreement levels are still achieved, showing the consistency of our annotation.
@@
Turk ASSET Newsela SimPA
Dimension GPT-4 Qwen Llama T5 GPT-4 Qwen Llama T5 GPT-4 Qwen Llama T5 Ref. GPT-4 Qwen Llama
Fluency $98.1$ $98.9$ $99.2$ $97.8$ $97.8$ $99.7$ $98.6$ $96.9$ $99.6$ $99.4$ $99.4$ $98.4$ $98.0$ $98.4$ $98.2$ $99.2$
Meaning $86.4$ $87.2$ $58.8$ $57.9$ $84.7$ $88.3$ $70.8$ $58.2$ $66.6$ $59.4$ $55.5$ $87.3$ $62.6$ $83.7$ $85.4$ $54.3$
Simplicity $77.7$ $90.5$ $82.2$ $85.5$ $68.2$ $83.0$ $64.1$ $90.3$ $81.3$ $91.2$ $82.4$ $95.5$ $76.5$ $88.5$ $91.7$ $90.2$
@@ 6. Annotation Result Analysis
Our comprehensive analysis of annotation data reveals that, overall, LLMs generate fewer erroneous simplification outputs compared to Control-T5, demonstrating higher ability in simplification. Larger models, such as GPT-4 and Qwen2.5-72B, excel at preserving meaning compared to smaller models like Llama-3.2-3B and Control-T5. However, larger LLMs are not without flaws; their most common error is replacing simpler lexical expressions with more complex ones.
@@ 6.1. Analysis of Task 1: Error Identification
This section presents a comparative analysis of erroneous simplification outputs generated by GPT-4, Qwen2.5-72B, Llama-3.2-3B, and Control-T5, focusing on error quantification and type analysis. We assess erroneous simplification outputs across four datasets: Turk, ASSET, Newsela, and SimPA, defining an erroneous output as one containing at least one error. As mentioned in Section 4.2, we did not include SimPA for Control-T5. To reduce bias stemming from this particular dataset, we also report the results after excluding SimPA, i.e., only on Turk, ASSET, and Newsela (denoted as ‘T&A&N’). Figure 3 shows that the three LLMs generally with fewer errors than Control-T5. This performance difference underscores LLMs’ superior performance in simplification tasks. Within the LLM group, Qwen2.5-72B produced the fewest errors, followed by GPT-4, while Llama-3.2-3B generated the most.
Figure 3. Comparison of the number of erroneous simplification outputs generated by models. Numbers in the square brackets represent the number of samples in datasets.